# Open-World Evaluation for Retrieving Diverse Perspectives

### v1

@@ Abstract
We study retrieving a set of documents that covers various perspectives on a complex and contentious question (e.g., will ChatGPT do more harm than good?). We curate a Benchmark for Retrieval Diversity for Subjective questions (BeRDS), where each example consists of a question and diverse perspectives associated with the question, sourced from survey questions and debate websites. On this data, retrievers paired with a corpus are evaluated to surface a document set that contains diverse perspectives. Our framing diverges from most retrieval tasks in that document relevancy cannot be decided by simple string matches to references. Instead, we build a language model based automatic evaluator that decides whether each retrieved document contains a perspective. This allows us to evaluate the performance of three different types of corpus (Wikipedia, web snapshot, and corpus constructed on the fly with retrieved pages from the search engine) paired with retrievers. Retrieving diverse documents remains challenging, with the outputs from existing retrievers covering all perspectives on only 33.74% of the examples. We further study the impact of query expansion and diversity-focused reranking approaches and analyze retriever sycophancy. Together, we lay the foundation for future studies in retrieval diversity handling complex queries.
@@ 1 Introduction
Figure 1: Overview of our benchmark and task. The upper part shows example instances. Each instance consists of a question and a perspective set. The lower part illustrates the task. The retriever will return a set of documents, and we evaluate whether the retrieved set (in this example, consisting of two documents) contains multiple answers or perspectives. MRecall measures the coverage of answers and perspectives, and precision measures whether each document contains any perspectives. A separate module (denoted as “Perspective Detection”) is developed to determine whether a document contains a certain perspective. We experiment with retrieving from Web corpora instead of Wikipedia, which provides insufficient information for the task.
Given a complex and contentious question Xu et al. (2024), such as “Will ChatGPT do more harm than good?”, a retrieval system should be able to surface diverse opinions in their top retrieval outputs. However, existing information retrieval (IR) tasks and systems mostly optimize for relevance to the question, ignoring diversity. Surfacing diverse documents can be useful to the users directly, but also can be improve retrieval-augmented language models (RALMs). Prompting large language models (LLMs) to generate an answer that encompasses diverse perspectives on its own is challenging Sorensen et al. (2024); Hayati et al. (2023a), and retrieval-augmentation Divekar and Durrett (2024) can facilitate LLMs to generate more comprehensive answers that represent diverse perspectives.
Therefore, we study retrieval diversity, evaluating whether a set of top-scored outputs from the retrieval system cover diverse perspectives for an input query. We loosely define perspective as a particular attitude or viewpoint towards the question expressed towards the input query.  Figure 1 shows examples of possible perspectives for different input questions. To tackle this task, we curate a Benchmark for Retrieval Diversity for Subjective questions, BeRDS, which contains 3K complex questions, each paired with on average 2.3 perspectives. The questions come from three sources: survey question collection Santurkar et al. (2023), debate question corpus Wachsmuth et al. (2018), and newly scrapped data from a debate platform (Kialo) with argument maps.
@@
Having constructed an evaluation framework, we test retrievers Robertson et al. (2009); Karpukhin et al. (2020a); Izacard et al. (2022) and different corpus. Given the subjectivity of the questions we consider, Wikipedia alone is insufficient for answering many questions in our proposed datasets. Therefore, we use Sphere Piktus et al. (2021), a subset of a web snapshot (CCNet Wenzek et al. (2020)), and a retrieval corpus built on the fly for each question from the Google search engine output. We find that pairing a dense retriever with web corpora (Sphere and Google Search) yields more diverse outputs compared to other settings. Yet, our experimental result suggests that current retrievers surface relevant documents but cannot present document sets with diverse perspectives, even when retrieving from the web.
To enhance the diversity of the retrieval results, we implement simple re-ranking Carbonell and Goldstein (1998) and query expansion approaches Mao et al. (2021). The former re-computes the scores for each document by penalizing documents that are similar to the documents that were retrieved before. The latter first generates multiple perspectives regarding the question using LLM (GPT-4) and uses them to guide diverse retrieval. Encouragingly, both methods showed strong gains on dense base retriever Izacard et al. (2022). We further provide rich analysis studying the coverage of each corpus, retriever sycophancy, and whether retrievers prefer supporting or opposing perspectives to the input query.
All our code, data, and trained model checkpoints are released at https://timchen0618.github.io/berds/ to promote future work in studying retrieval diversity for complex questions.
@@ 2 Related Work
Human society possesses pluralistic values Sorensen et al. (2024) and exhibits diverse perspectives in various scenarios, including responses to survey questions Santurkar et al. (2023); Durmus et al. (2023), news editorials Liu et al. (2021), and stance with respect to debate or controversial questions Chen et al. (2019, 2022b); Wan et al. (2024). Hayati et al. (2023b) looks into eliciting diverse perspectives from LLM on social norm topics. Wan et al. (2024) studies LLM generation when in-context documents provide contradicting statements. Concurrent work Xu et al. (2024) studies evaluating perspective diversity of debatable questions. One way to encourage diversity in LLM output is retrieval augmentation, by providing documents with diverse perspectives. We focus on diverse retrieval instead of generation from LLMs.
@@
Previous works have investigated retrieval diversity in the context of entity ambiguity Clarke et al. (2008); Agrawal et al. (2009); Wang et al. (2017). Closer to our definition of diversity,  Chen et al. (2019) propose a task substantiated perspective discovery, where systems are supposed to discover a diverse set of perspectives taking stance against a given claim.  Chen et al. (2022b) further argues for the need of a multi-perspective search engine that provides more direct and diverse responses. Despite these efforts, there is no existing benchmark for retrieving diverse perspectives in realistic corpora. To facilitate progress towards retrieving diverse perspectives for subjective questions, we formalize the task and construct a benchmark for evaluation, paired with automatic evaluation metrics. Concurrent work Ziems et al. (2024) studies the bias of retrieval systems along a single polar axis (e.g. capitalism vs. socialism) for debatable questions and develops an unsupervised automatic metric.
Concurrent work Zhao et al. (2024) explores the perspective-awareness of retrieval systems, where they repurpose existing datasets to build a retrieval benchmark where queries are supplemented with a perspective. While sharing a similar goal, they evaluate retrieval with perspectives as input and measure if systems follow them, while in our settings the retrievers are not given target perspectives. Furthermore, our work expands knowledge sources significantly using a web corpus, while their work focuses on much smaller corpora.
@@ 3 Task Formulation
Min et al. (2021b) extended retrieval task to cover questions that have multiple valid answers. Here, each question $q$ is paired with gold answer set $\{a_{1},a_{2},...a_{m}\}$ where each gold answer $a_{i}$ is a short string. The goal is to retrieve a diverse set of documents such that the retrieval output will contain all $m$ answers. New evaluation metric MRecall @ $k$ is defined as (1) when number of answers $m>k$, MRecall @ $k$ = 1 if the retrieved document set $D$ contains all $k$ answers. (2) when $m\leq k$, MRecall @ $k$ = 1 if the retrieved document set $D$ contains all $m$ answers. Otherwise MRecall @ $k$ = 0. Simple string matching determines if a document $d_{i}$ contains answer $a$.
@@
Table 1: Dataset statistics. The second column denotes the number of examples per dataset. We spare 25% of each dataset as the development (Dev.) set, and use the remainder as the test set. # $p$ denotes the average number of perspectives per question, and $|p|$ denotes the average number of words (by NLTK word_tokenize) in each perspective.
@@
We consider questions on subjective topics that could be approached from different perspectives, and there is no fixed or complete set of reference answers. We assume a set of $m$ perspectives $\{p_{1},p_{2}...,p_{m}\}$ for addressing question $q$. The goal of a retriever is to output a diverse set of documents covering all $m$ perspectives.
@@
Deciding if a document contains an answer can be approximated by simple string matching, but deciding if a document supports a perspective is nontrivial. The most accurate metric would be collecting human judgments. However, this is not scalable, so we build a model that approximates human judgment (presented in Appendix A). We refer to this task of determining if a document contains a certain perspective as “Perspective Detection”. We do not use off-the-shelf stance detection models Zhang et al. (2023a); Alturayeif et al. (2023) as the perspectives might not be binary. A single document could contain more than one perspective.
@@
We use MRecall @ $k$ from prior work Min et al. (2021b). The key difference is instead of checking whether a document contains any of answer $\{a_{1},a_{2},...a_{m}\}$, we check whether a document supports any of perspectives $\{p_{1},p_{2}...,p_{m}\}$. We also report Precision @ $k$, the percentage of documents in the retrieved document set $D$ that contain any perspectives. This measures the relevance of the retrieved documents to the question.
This evaluation setup does not assume annotated “gold” documents or answers for each question, enabling comparing different corpora. We consider the corpus in Section 4.2 and discuss the effects of corpora in Section 7.
@@ 4 The BeRDS Benchmark
The BeRDS Benchmark aims to measure retrieval diversity for questions that are opinionated or invite diverse perspectives. A single instance in BeRDS will consist of a question $q$ and a set of $m$ perspectives $\{p_{1},p_{2}...,p_{m}\}$. Each perspective should be relevant to the question and sufficiently different from each other. We provide the data selection and construction process below.
@@ 4.1 Data Construction
We aim to look for questions that elicit diverse valid perspectives naturally. Questions related to debate topics would be good candidates since they would induce at least two contradicting, valid viewpoints. Another natural choice is survey questions, where people would express differing opinions. We thus collect questions from a debate website (kialo.com) and a dataset of survey questions, OpinionQA Santurkar et al. (2023). We also generate questions from Arguana Wachsmuth et al. (2018), a dataset of contradicting arguments. We provide examples for each dataset in BeRDS in Figure 1, and data statistics in Table 1. For both Arguana and OpinionQA, we only consider two contradictory perspectives, one supporting and one opposing. These perspectives may not be exhaustive but cover a range of arguments for the question. We additionally provide examples of the raw data and generated data in Table 11.
@@
Table 2: Retrieval corpus statistics. # pas. denotes the number of passages; # doc. denotes the number of documents, where each document is a 100-word passage. For Google Search output, we construct a corpus for each example, and we report the average statistics across examples.
@@
is a dataset of survey questions on various topics including privacy, political views, and health, targeted toward the citizens of the United States. The dataset originates from the annual Pew American Trends Panel survey. The dataset contains a wide range of survey questions, such as those asking about personal experiences (Regardless of whether or not you own a gun, have you ever fired a gun?). We filter such questions by prompting GPT-4, removing 22.7% of the questions (345 out of 1521 in total). The survey questions come with multiple options, with each representing a certain degree of support for the question. Conditioned on the survey question and the options, we convert each survey question into a natural language question and generate one supporting and one opposing perspective with GPT-4.
We use gpt-4-0613 for all the data generation. All the input prompts for generation can be found in Section B.1. The total cost for generating the data is about \$100 USD.
@@ 4.2 Retrieval Corpus
Prior works on single-answer or multi-answer retrieval usually adopt Wikipedia as the main retrieval corpus. However, Wikipedia alone cannot cover all the information required Redi et al. (2020). This is even more likely as our questions are inherently opinionated and Wikipedia texts are supposed to be neutral. We propose to use the Sphere Piktus et al. (2021) corpus, which is a subset of CCNet Wenzek et al. (2020), as the knowledge source. One advantage of retrieving from web corpus over using a commercial search engine (e.g. Google Search) results is reproducibility.
@@
Instead of building a static corpus and retrieving documents from it, one can use a web search engine to retrieve new documents on the fly. This enables the system to access bigger, more up-to-date source documents. The downside of using web search API  Nakano et al. (2021); Yoran et al. (2023) is the lack of reproducibility as well as potentially returning documents from less credible sources. [^1] Prior work Chen et al. (2024) studied the reproducibility of search API results, comparing URLs between the sets of documents retrieved at different timestamps with the same query. They report only 30% of URLs overlap when queried two months apart.
We implement web search by retrieving documents from Google Search API. [^2] https://serper.dev/ We first obtain the raw HTML of top hundred documents, which are then processed and converted to plain text using html2text [^3] https://pypi.org/project/html2text/ and readability, [^4] https://github.com/buriy/python-readability following Chen et al. (2022a, 2024). We split the processed documents into 100-word segments and take all the segments as the corpus. Unlike Wikipedia and Sphere, we build a separate corpus for each question.
Corpus Model Datasets Average
@@
Google Search BM25 35.15 21.07 35.43 22.61 41.27 21.09 37.28 21.59 43.02
Output Contriever 52.32 27.6 50.31 28.94 48.91 22.79 50.51 26.44 57.56
Table 3: Performance of base retrievers on the test split of BeRDS. In the last three columns, we report macro-average over the three datasets.
5 Retrieval Approaches
We consider a suite of standard retrievers (sparse and dense) on our dataset. As these retrievers do not encourage the diversity of their output, we implement a re-ranking method and a query expansion method.
Baselines
We implement BM25 Robertson et al. (2009), a widely adopted bag-of-word retrieval function considering token-matching between questions and documents. We use the LuceneSearcher of Pyserini package to build the Wikipedia corpus. For the Sphere corpus, we use the faiss index provided by  Piktus et al. (2021).
We consider two dense retrievers, DPR Karpukhin et al. (2020a) and Contriever Izacard et al. (2022). The former is a dual encoder trained with contrastive learning on the Natural Questions (NQ) dataset Kwiatkowski et al. (2019). The latter is an encoder trained with a self-supervised contrastive objective. DPR uses separate encoders for document and question embeddings, and Contriever uses a shared encoder. Both models are initialized from a BERT-base encoder. We use the version of Contriever further fine-tuned on the MSMARCO dataset Nguyen et al. (2016) after self-supervised learning, which performed better in the original paper.
Re-ranking
where $R$ denotes a set of top retrieved documents considered for re-ranking ($|R|=100$ in our experiments), $S$ denotes the document set that is already selected, and $\lambda$ is a hyperparameter. We tune $\lambda$ on the development split. [^5] We normalize the $Sim_{1}$ scores to be between 0 and 1, and consider $\lambda$ values of [0.5, 0.75, 0.9, 0.95, 0.99]. $Sim_{1}$ is the score from the retriever itself. To compute $Sim_{2}$, we implement the Universal AnglE Embedding Li and Li (2023) to obtain the document embeddings and take the cosine similarity between two documents. We choose this embedding model as it is the top-ranked open model on the MTEB leaderboard Muennighoff et al. (2023) with less than 1B parameters (335M). A more compact model is preferable as we compute document embeddings for the top hundred documents per example. The experimental details can be found in Section B.6.
Query Expansion
[…]
Corpus Retriever Default + Re-ranking + Query Expansion
Prec. MRec. Prec. MRec. Prec. MRec.
DPR 17.33 6.21 16.62 (-4.1%) 6.06 (-2.5%) 17.78 (+2.6%) 6.65 (+7.1%)
Table 4: Performances on BeRDS after re-ranking and diversity-aware retrieval. Each cell reports the macro-average over three datasets in BeRDS. Bolded numbers are the best performance in each row. Blue numbers indicate increases compared to the default setting.
6 Experimental Results
We report MRecall @ $5$ and Precision @ $5$, [^7] We additionally report the results for $k$ = 10 in Section C.1, where very similar trends hold. as defined in Section 3. All evaluation metrics require judging whether each retrieved document contains a perspective (defined as perspective detection, and more details can be found in Appendix A). We use a fine-tuned Mistral-7B model as the evaluator. [^8] We will open-source this model to promote future work.
Table 3 presents the performance of base retrievers on BeRDS using three retrieval corpus described in Section 4.2. We do not report DPR results on Google Search outputs as we find it does not generalize well beyond the Wikipedia corpus.
Effect of Corpus.
Comparing retrievers operating on Wikipedia and Sphere, most retrievers achieve higher MRecall and Precision when retrieving from Sphere, except DPR. We hypothesize that DPR is trained only on the documents from the Wikipedia corpus so that it is not generalizing to the web corpus. Taking web search results (Google Search) as a corpus shows improved diversity over Wikipedia, but lags behind retrieving from the entire web (Sphere). This might be caused by the gap in the corpus size – we only take the top 100 Google Search results and build corpora with them (resulting in a corpus size of 4.4K passages on average), while Sphere contains 906M passages.
Comparing Retrievers.
Across all experimental settings, Contriever achieves the highest diversity, followed by BM25 and DPR. BM25 generalizes better than DPR, which is fine-tuned on NQ with Wikipedia corpus. Even the best setting (Contriever on Sphere) only achieves an average MRecall of 30.64. This indicates even the best retriever struggles to retrieve a comprehensive set of documents. Precision @ 1 is as high as 58.67, showing that retrievers could surface one of the perspectives effectively.
Results: Approaches to Improve Diversity
We report the results of re-ranking and query expansion approaches in Table 4. Re-ranking improves diversity on four of eight retriever-corpus settings, at the cost of Precision for all settings. The limited performance gains could be because semantic dissimilarity does not always infer differing perspectives. It is encouraging that we observe gain in Contriever settings, our best retriever. Query expansion improves the Precision for five settings and boosts MRecall for six settings (over 10% for three of them). This indicates that GPT-4 could diversify the search space by providing queries hinting different perspectives to the retrievers. Future work can explore combining both approaches.
7 Analysis
Corpus Model Avg. size ($\downarrow$) % within top 100 ($\uparrow$)
Wiki BM25 29.8 41.5
DPR 30.9 36.7
Cont. 24.7 55.6
Sphere BM25 20.8 77.3
DPR 37.4 48.9
Cont. 17.8 83.1
Google BM25 21.0 72.1
Cont. 18.2 76.2
How many documents should one retrieve to cover diverse perspectives?
So far, we have reported the performances of the retrieval outputs consisting of five documents and found that the model outputs do not cover diverse perspectives. If we consider up to the top 100 documents, can we cover diverse perspectives? What is the average number of documents needed to be retrieved to cover all perspectives? Table 5 reports the size of the retrieved document set needed to cover all the perspectives with base retrievers. We find that the best setting (Contriver with Sphere) can find all the perspectives within the top 100 document set 83.1% of the time, and for those examples, the average number of documents required to cover all the perspectives is 17.8. Other settings show worse performance, mostly reflecting performance reported in Table 3.
Retrievers often fail to surface documents that cover comprehensive perspectives. Can we characterize the perspective that is more likely to be surfaced? We focus on questions with only two perspectives, one taking the supporting stance towards the question and the other opposing. [^9] We identify this by prompting GPT-4; details and prompts can be found in Section B.8. We report whether the top five documents obtained by the retrievers cover only supporting, opposing, both, or neither perspectives in Figure 2. We observe that all retrievers retrieve the supporting perspectives significantly more often when they fail to retrieve both.
BM.DPRCo.BM.DPRCo.BM.Co.000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Wiki.               Sphere           Googlebothnonesupportingopposing
Figure 2: We report whether retrievers prefer the supporting or the opposing perspectives. Results are computed on the top-5 documents aggregated over three datasets. All retrievers favor the supporting perspectives. BM. denotes BM25 retriever, Co. denotes Contriever, and Wiki. denotes the Wikipedia corpus.
Retriever Sycophancy.
We quantify leaning towards supporting perspectives to be $\Delta$ = ($p$ $n$)/$p$, where $p$ and $n$ are the percentages of documents containing supporting and opposing perspectives, respectively. Greater $\Delta$ indicates the retrievers favor the supporting perspectives more. We present the results on the Sphere corpus in Figure 3.101010Results on the other corpora show the same trends, except for DPR on the web corpus. Querying the retrievers with supporting perspectives increases $\Delta$ across the board, and retrieving with opposing perspectives decreases $\Delta$. Retrievers tend to favor perspectives that they are prompted with. This partially explains the gains from the query expansion.
001010101020202020303030304040404050505050Wiki_BM25Wiki_DPRWiki_Cont.supporting perspectivedefault queryopposing perspective
Is retrieval diversity limited by the corpora?
Unlike prior work, we do not assume a gold corpus that contains target information. The best MRecall @ $5$ on BeRDS achieved by any approach on average is below 35. How much of the failure is caused by the limitation of the corpus or that of retrievers? We answer this question by computing MRecall over a larger document set, where we combine the top 100 retrieved documents from all retrievers for each corpus. If this number nears 100, we could theoretically find an optimal subset of $k$ documents for each question that achieve MRecall @ $k$ close to 1. We present the results in Table 6. Only the Wikipedia corpus contain insufficiently diverse information, achieving MRecall of 62% when combining three retriever outputs. In contrast, the web corpora do not limit retrievers’ performances in diversity. This result hints at great room for improvement for the retrievers.
Corpus Arguana Kialo OpinionQA Average
Wiki. 71.87 61.24 55.10 62.74
Sphere 92.80 85.27 89.46 89.18
Google 82.67 82.69 77.66 81.01
Table 6: MRecall over the union of the top 100 retrieved documents from all retrievers for each corpus. These values are substantially higher than MRecall @ $5$, suggesting the corpora (especially Sphere) have good information coverage.
8 Conclusion
We study retrieval diversity for complex, subjective questions. We curate a benchmark BeRDS that evaluates retrieval diversity, provides automatic metrics, and reports the performances of several retriever baselines. Our experiments show that current retrievers cannot retrieve comprehensive information in the proposed setting.
Once we retrieve a set of diverse documents, summarizing such rich information into a coherent response Laban et al. (2022); Huang et al. (2023); Zhang et al. (2023b) could be useful for the users. Connecting our work, which enables uncovering diverse documents, with summarization can be a fruitful direction for future work.
Limitations
  • The perspective set in our dataset is mostly binary (the average number of perspectives is 2.9 for Kialo and 2.0 for Arguana and OpinionQA). More fine-grained perspectives could be explored.
Acknowledgements
We thank the members of UT NLP, especially Fangyuan Xu, for their feedback on the draft. The project is partially supported by a grant from Open Philanthropy and NSF grant RI-2312948. We also thank Chitrank Gupta for contributing to the earlier phase of this work.
Appendix A Identifying whether a perspective is represented in a document

### v2

@@ Abstract
We study retrieving a set of documents that covers various perspectives on a complex and contentious question (e.g., will ChatGPT do more harm than good?). We curate a Benchmark for Retrieval Diversity for Subjective questions (BeRDS), where each example consists of a question and diverse perspectives associated with the question, sourced from survey questions and debate websites. On this data, retrievers paired with a corpus are evaluated to surface a document set that contains diverse perspectives. Our framing diverges from most retrieval tasks in that document relevancy cannot be decided by simple string matches to references. Instead, we build a language model-based automatic evaluator that decides whether each retrieved document contains a perspective. This allows us to evaluate the performance of three different types of corpus (Wikipedia, web snapshot, and corpus constructed on the fly with retrieved pages from the search engine) paired with retrievers. Retrieving diverse documents remains challenging, with the outputs from existing retrievers covering all perspectives on only 40% of the examples. We further study the effectiveness of query expansion and diversity-focused reranking approaches and analyze retriever sycophancy.
@@ 1 Introduction
Figure 1: Overview of our benchmark and task. The upper part shows example instances, where each instance consists of a question and a perspective set. The lower part illustrates the task. The retriever will return a set of documents, and we evaluate whether the retrieved set (in this example, consisting of two documents) contains multiple answers or perspectives. MRecall measures the coverage of answers and perspectives, and precision measures whether each document contains any perspectives. A separate module (denoted as “Perspective Detection” (Section 5)) is developed to determine whether a document contains a certain perspective. We experiment with retrieving from the Web in addition to Wikipedia, which provides insufficient information for the task.
Given a complex and contentious question Xu et al. (2024), such as “Will ChatGPT do more harm than good?”, a retrieval system should be able to surface diverse opinions in their top retrieval outputs. However, existing information retrieval (IR) tasks and systems mostly optimize for relevance to the question, ignoring diversity. A diverse IR system could be useful in two ways: (1) by surfacing diverse documents, which can be useful to users directly Chen et al. (2022b), and (2) by improving retrieval-augmented language models (RALMs). Prompting large language models (LLMs) to generate an answer that encompasses diverse perspectives on its own is challenging Sorensen et al. (2024); Hayati et al. (2023a), and retrieval-augmentation Divekar and Durrett (2024) can facilitate LLMs to generate more comprehensive answers that represent diverse perspectives. Yet, existing retrieval benchmarks do not focus on evaluating retrieval diversity.
In this work, we study retrieval diversity, building a framework for evaluating whether a set of top-scored outputs from the retrieval system cover diverse perspectives [^1] We loosely define perspective as a particular attitude or viewpoint expressed towards the input question. for an input query.  Figure 1 shows examples of possible perspectives for different input questions. To tackle this task, we curate a Benchmark for Retrieval Diversity for Subjective questions, BeRDS, which contains 3K complex questions, each paired with on average 2.3 perspectives. The questions come from three sources: survey question collection Santurkar et al. (2023), debate question corpus Wachsmuth et al. (2018), and newly scrapped data from a debate platform (Kialo) with argument maps.
@@
Having constructed an evaluation framework, we test popular retrievers Robertson et al. (2009); Karpukhin et al. (2020a); Izacard et al. (2022) and three corpora. Given the subjectivity of the questions we consider, Wikipedia alone is insufficient for answering many questions. Therefore, we use Sphere Piktus et al. (2021), a subset of a web snapshot (CCNet Wenzek et al. (2020)), and a retrieval corpus built on the fly for each question from the Google search engine output. We find that pairing a dense retriever with web corpora (Sphere and Google Search) yields the most diverse outputs. Yet, our experimental result suggests that current retrievers surface relevant documents but cannot present document sets with diverse perspectives, even when retrieving from a richer web corpus.
To enhance the diversity of the retrieval results, we implement simple re-ranking Carbonell and Goldstein (1998) and query expansion approaches Mao et al. (2021). The former re-computes the scores for each document by penalizing documents that are similar to the documents that were retrieved before. The latter first generates multiple perspectives regarding the question using LLM (GPT-4) and uses them to guide diverse retrieval. Encouragingly, both methods showed gains on the strongest dense base retriever Izacard et al. (2022), despite not boosting performances for all settings. We further provide rich analysis studying the coverage of each corpus, retriever sycophancy, and whether retrievers prefer supporting or opposing perspectives to the input query.
All our code, data, and trained model checkpoints are released to promote future work in studying retrieval diversity for complex questions. [^2] https://timchen0618.github.io/berds/
@@ 2 Related Work
Human society possesses pluralistic values Sorensen et al. (2024) and exhibits diverse perspectives in various scenarios, including responses to survey questions Santurkar et al. (2023); Durmus et al. (2023), news editorials Liu et al. (2021), and stance with respect to debate or controversial questions Chen et al. (2019, 2022b). Hayati et al. (2023b) looks into eliciting diverse perspectives from LLM on social norm topics. Wan et al. (2024) studies LLM generation when in-context documents provide contradicting statements. Concurrent work Xu et al. (2024) studies evaluating perspective diversity of debatable questions. One way to encourage diversity in LLM output is retrieval augmentation, by providing documents with diverse perspectives. We focus on retrieving diverse opinions from existing corpus instaed of generating diverse opinions from LLMs.
@@
Previous works have investigated retrieval diversity in terms of entity ambiguity Clarke et al. (2008); Agrawal et al. (2009); Wang et al. (2017). Closer to our definition of diversity,  Chen et al. (2019) propose a task substantiated perspective discovery, where systems are supposed to discover diverse perspectives taking stances against a given claim.  Chen et al. (2022b) further argues for the need of a multi-perspective search engine that provides direct and diverse responses. Despite these efforts, there is no benchmark for retrieving diverse perspectives in realistic corpora. We formalize such task and construct evaluation benchmark and metrics. Concurrent work Ziems et al. (2024) studies the bias of retrievers along a polar axis (e.g. capitalism vs. socialism) for debatable questions and develops an unsupervised automatic metric.
Concurrent work Zhao et al. (2024) explores the perspective-awareness of retrieval systems, where they repurpose existing datasets to build a retrieval benchmark where queries are supplemented with a perspective. While sharing a similar goal, they evaluate retrieval with perspectives as input and measure if systems follow them, and in our settings, the retrievers are not given target perspectives. Furthermore, our work expands knowledge sources significantly using a web corpus, while their work focuses on much smaller corpora.
@@ 3 Task Formulation
Min et al. (2021b) extended retrieval task to cover questions with multiple valid answers. Here, each question $q$ is paired with gold answer set $\{a_{1},...a_{m}\}$ where each answer $a_{i}$ is a short string. The goal is to retrieve a diverse set of documents such that the retrieval output will contain all $m$ answers. New evaluation metric MRecall @ $k$ is defined as (1) when number of answers $m>k$, MRecall @ $k$ = 1 if the retrieved document set $D$ contains $k$ answers. (2) when $m\leq k$, MRecall @ $k$ = 1 if $D$ contains all $m$ answers. Otherwise MRecall @ $k$ = 0. Simple string matching determines if a document $d_{i}$ contains answer $a$.
@@
Table 1: Dataset statistics. The second column denotes the number of examples. We spare 25% of each dataset as the development (Dev.) set, and use the remainder as the test set. # $p$ is the average number of perspectives per question, and $|p|$ denotes the average number of words (by NLTK word_tokenize) in each perspective.
@@
We consider questions on subjective topics that could be approached from different perspectives, without fixed or complete sets of reference answers. We assume a set of $m$ perspectives $\{p_{1},...,p_{m}\}$ for addressing question $q$. Retrievers should output a diverse set of documents covering all $m$ perspectives.
@@
Deciding if a document contains an answer can be approximated by simple string matching, but deciding if a document supports a perspective is nontrivial. The most accurate metric would be human judgments. However, this is not scalable, so we build a model approximating human judgment. We refer to this task of determining if a document contains a certain perspective as “Perspective Detection” (presented in Section 5). We do not use off-the-shelf stance detection models Zhang et al. (2023a); Alturayeif et al. (2023) as the perspectives might not be binary. A single document could contain more than one perspective.
@@
We use MRecall @ $k$ from prior work Min et al. (2021b). The key difference is instead of checking whether a document contains any of answer $\{a_{1},...,a_{m}\}$, we check whether a document supports any of perspectives $\{p_{1},...,p_{m}\}$. We also report Precision @ $k$, the percentage of documents in the retrieved document set $D$ that contain any perspectives. This measures the relevance of $D$ to the question.
This evaluation setup does not assume annotated “gold” documents or answers for each question, enabling comparing different corpora. We discuss the effects of corpora in Section 8.
@@ 4 The BeRDS Benchmark
The BeRDS Benchmark aims to measure retrieval diversity for questions that are opinionated or invite diverse perspectives. A single instance in BeRDS will consist of a question $q$ and a set of $m$ perspectives $\{p_{1},p_{2}...,p_{m}\}$. Each perspective should be relevant to the question and distinct. We provide the data selection and construction process below.
@@ 4.1 Data Construction
We aim to look for questions that elicit diverse valid perspectives naturally. Questions related to debate topics would be good candidates since they would induce at least two contradicting, valid viewpoints. Another natural choice is survey questions, where people would express differing opinions. We thus collect questions from a debate website (kialo.com) and a dataset of survey questions, OpinionQA Santurkar et al. (2023). We also generate questions from Arguana Wachsmuth et al. (2018), a dataset of contradicting arguments. We provide examples for each dataset in BeRDS in Figure 1, and data statistics in Table 1. For both Arguana and OpinionQA, we only consider two contradictory perspectives, one supporting and one opposing. These perspectives may not be exhaustive but cover a range of arguments for the question.  [^3] We mostly focus on binary perspectives since they are well-defined for contentious queries. Future work could explore more fine-grained and multi-dimensional perspectives. We provide examples of the raw data and generated data in Table 10.
@@
Table 2: Retrieval corpus statistics. # pas. is the number of passages; # doc. is the number of documents, where each document is a 100-word passage. For Google Search output, we construct a corpus for each example, and we report the average statistics across examples.
@@
is a dataset of survey questions on various topics including privacy, political views, and health, targeted toward the citizens of the United States. The dataset originates from the annual Pew American Trends Panel survey. The dataset contains questions asking about personal experiences (Regardless of whether or not you own a gun, have you ever fired a gun?), which we filter by prompting GPT-4, removing 22.7% of the questions (345 out of 1521 in total). The survey questions come with multiple options, with each representing a certain degree of support for the question. Conditioned on the survey question and the options, we convert each survey question into a natural language question and generate one supporting and one opposing perspective with GPT-4.
We use gpt-4-0613 for all data generation. All the input prompts for generation can be found in Section A.1. The total cost is about \$100 USD. We manually examine 20 examples of each subset, and only two examples (out of 60) contain minor errors (see Appendix B.4 for examples).
@@ 4.2 Retrieval Corpus
Prior works on retrieval usually adopt Wikipedia as the main retrieval corpus. However, Wikipedia alone has limited knowledge coverage Redi et al. (2020). This is even more likely in our case of complex and subjective queries, as Wikipedia is supposed to be neutral. One downside of moving from Wikipedia to a Web snapshot is potentially returning documents from less credible sources. We use the Sphere Piktus et al. (2021) corpus, which is a subset of CCNet Wenzek et al. (2020), as the web knowledge source.
@@
Instead of building a static corpus and retrieving documents from it, one can use a web search engine to retrieve new documents on the fly. This enables the system to access bigger, more up-to-date source documents. The downside of using web search API  Nakano et al. (2021); Yoran et al. (2023) is the lack of reproducibility. [^4] Prior work Chen et al. (2024) studied the reproducibility of search API results, comparing URLs between the sets of documents retrieved at different timestamps with the same query. They report only 30% of URLs overlap when queried two months apart.
We will use Google Search API to build web corpus on the fly. [^5] https://serper.dev/ We first obtain the raw HTML of top hundred documents, which are then processed and converted to plain text using html2text [^6] https://pypi.org/project/html2text/ and readability, [^7] https://github.com/buriy/python-readability following Chen et al. (2022a, 2024). We split the processed documents into 100-word segments and take all the segments as the corpus. Unlike Wikipedia and Sphere, we build a separate corpus for each question.
Corpus Model Datasets Macro-Average
@@
Tart 65.47 43.60 56.47 32.80 63.25 31.60 61.73 36.00 65.16
NV-Embed-v2 69.84 43.87 66.85 40.31 64.94 34.35 67.21 39.51 71.27
BM25 35.15 21.07 35.43 22.61 41.27 21.09 37.28 21.59 43.02
Google Search Contriever 52.32 27.6 50.31 28.94 48.91 22.79 50.51 26.44 57.56
Output Tart 53.33 30.13 50.48 27.87 50.83 24.67 51.55 27.56 59.07
NV-Embed-v2 62.40 39.87 64.19 37.60 56.88 28.67 61.16 35.38 67.87
Table 3: Performance of retrievers on the test split of BeRDS. We do not report NV-Embed-v2 results on Sphere for computational limitations. Overall, retrievers show better performance with Sphere corpus, and NV-Embed-v2 shows the strongest performance in both corpora that it was evaluated on.
5 Perspective Detection